OpenStudy (anonymous):

how to find the focus and directrix for the parabola y=(x/4)^2 (squared)

OpenStudy (anonymous):

Change it into standard form: \[x^2 = 4ay\] \[y = x^2/16 \] \[x^2=16y\] Find a: 4a=16 a=4 And then you find the vertex, which is (0,0) From there you can deduce that the focus will be (0,4) Directrix is y=-4

OpenStudy (anonymous):

good